3x+3y-3z=-9
y+4z=-29
2x-3y. =7
I don't know how to eliminate this one because the x is already gone in one and the z is gone in the other one.

Answer by richwmiller(17219)

3x+3y-3z=-9
y+4z=-29
2x-3y =7
x=-4
y=-5
z=-6